What is the osmolality of a 0.030 M solution of lithium sulfate? 2*0.030+ 1*0.030=0.09 osmolality
Answer #1

Solution:

Osmolarity = \sum (molarity x coefficient of dissociation)

Lithium sulfate (Li2SO4) is dissociated as,

Li2SO4 === 2Li+ + SO4^2-

Coefficient of dissociation for Li+ = 2

Coefficient of dissociation for SO4^2- = 1

Hence,

Osmolarity = \sum (0.030 x 2 + 0.030 x 1)

Osmolarity = 0.060 + 0.030 = 0.090 Osmol/L